# Surrogate normal-forms for the numerical bifurcation and stability analysis of Navier-Stokes flows via machine learning

### v1

@@ Abstract
Inspired by the Equation-Free multiscale modelling approach, we show how the “embed-learn-lift” framework can be exploited for the construction of surrogate “normal-forms” (reduced-order models of minimal dimensionality) of high-fidelity Navier-Stokes simulations, that can in turn be used for computationally efficient and numerically accurate bifurcation and stability analysis tasks. The framework consists of four main steps. In the first step, we apply manifold learning to discover the intrinsic dimension of the emergent latent space of the complex, high-dimensional Navier-Stokes spatio-temporal dynamics over the parameter space. In the second stage, we construct “normal-forms” like reduced-order models (ROMs)-in the sense that we capture the correct dimension for describing the emergent dynamics- on the latent space using Gaussian Process regression (GPR). Then, based on the constructed surrogate “normal-form” models, which dramatically reduce computational complexity, we exploit the toolkit of numerical bifurcation analysis theory to construct the bifurcation diagrams in the latent spaces, and perform systematic stability analysis. This allows, for example, the continuation of branches of limit cycles emanating from Andronov-Hopf bifurcations; such an analysis is intractable for high-dimensional systems such as the Navier-Stokes equations using standard Jacobian-full based computations. Finally, by solving the pre-image problem, we reconstruct the bifurcation diagrams in the original high-dimensional space. The proposed methodology is demonstrated by focusing on two suitable test-bed flow configurations: the wake flow past an infinitely long circular cylinder and the planar sudden-expansion channel flow. These two flow systems exhibit Andronov-Hopf and pitchfork (symmetry-breaking) bifurcations by increasing the Reynolds number $Re$, respectively. The proposed methodology successfully identifies the appropriate dimensionality of the latent space of the Navier-Stokes equations—and based on them constructs ROMs referred to here as "surrogate normal-forms" using Gaussian Process regression (GPR)—that enable the accurate tracing of the bifurcating solutions and their stability analysis. This includes the accurate tracing of limit cycles, their period, and their stability characteristics, as determined through the computation of Floquet multipliers.
@@ 1 Introduction
On the other side, data-driven methods have been propelled in recent years by the exponential growth of machine learning (ML) and deep learning methodologies, which aim to construct surrogate ROMs from high-fidelity spatio-temporal simulations [28, 29, 30, 16]. The first step involves projecting the high-dimensional spatio-temporal flow dynamics onto a carefully chosen (low-dimensional) latent space. This in CFD is typically done via manifold learning techniques like Proper Orthogonal Decomposition (POD) [31, 32, 33, 34]. POD offers a direct, closed-form solution to the pre-image problem, i.e., the reconstruction of the CFD dynamics back in the original high-dimensional space. Nonlinear manifold learning algorithms, such as diffusion maps (DMs), have also been exploited  [35, 36, 37, 38, 39, 40, 41, 42] to deal with more complex dynamics, including chaotic flows [42]. In these approaches, in contrast to the straightforward closed-form mapping between full and reduced spaces provided by POD, the solution of the pre-image problem suffers from ill-posedness, and different techniques such as geometric harmonics and kernel ridge regression have been proposed (for a review, see [43, 44]). Machine learning-based architectures, such as autoencoders, have also been used for learning a set of coordinates that can parametrize the latent space dynamics [45, 46, 47, 48, 49]. On the part of the construction of surrogate ROMs via machine learning, various approaches have been proposed including Sparse identification of nonlinear dynamical systems (SINDy) [50, 51, 52, 53] coupled also with auto-encoders  [54, 55], feedforward/recurrent neural networks [56, 57, 58, 39, 59, 40, 60, 61, 62], cluster-based networks [63], spectral submanifolds [64] and gaussian process regression models (GPR) [65, 66, 67, 68]. For example, SINDy has been used to model a range of fluid flows, including laminar and turbulent wakes [51], convective [52] and shear [53] flows, and it has been also applied in combination with auto-encoders to learn latent space coordinates before finding the dynamical system to evolve [54]. Deep Neural Network-based Auto Encoders (DNN-AE) have been recently employed to identify reduced-order coordinates of a turbulent flow over a building-like structure on a nonlinear manifold [48]. Diffusion maps and neural networks have been leveraged to build predictive reduced-order models of a vertically falling liquid film flow in the case of sparse data [40]. A combination of $\beta$-variational autoencoders and transformers has been used to learn parsimonious and near-orthogonal reduced-order models for two-dimensional viscous flows in both periodic and chaotic regimes [69]. In [70, 71, 72, 73, 74], ROMs and bifurcation diagrams of equilibria are built from steady-state snapshots of Navier–Stokes simulations, targeting “stable bifurcating branches” using intrusive POD Galerkin projections with applications to triangular lid-driven cavity and sudden-expansion channel flows. This approach has also been used for the control of solutions on the steady-state unstable branch [73] and generalized in cases of fluid-structure interaction problems [74].
@@
Here, building on previous efforts rooted in the celebrated Equation-Free multiscale framework [75], we propose a four-stage full data-driven framework for the construction of surrogate “normal-forms” via machine learning that can approximate not only qualitatively (as the traditional normal-form models) but also quantitatively, with high accuracy, the emergent dynamics of the Navier-Stokes equations. At this point, we note that we use the term “normal-forms” to emphasize the (main) challenge in complex systems modelling, that of learning the correct dimensionality of the latent dynamics across the parameter space, rather than reproducing the specific ROMs with minimal polynomial structures or local truncations as meant in local bifurcation theory. Consequently, our “normal-forms” for the Navier-Stokes equations represent ROMs with the correct dimensionality needed to approximate quantitatively the dynamics beyond the limited vicinity of the bifurcation points, which classical normal-forms cannot. Based on the fully data-driven constructed surrogate models, we then exploit the power of state-of-the-art continuation toolboxes such as MATCONT [3, 76] and AUTO [2], to perform computationally efficient and highly accurate bifurcation and stability analysis of the Navier-Stokes equations. Hence, we dramatically reduce computational and memory demands, yet still allow all the state-of-the-art bifurcation and stability (continuation of solution branches, even of unstable steady-states, continuation of limit cycles and their exact period, branch switching, the calculation of eigenvalues and Floquet multipliers, continuation of codimension-two bifurcations, continuation of homoclinic and heteroclinic bifurcations, Bautin and Takens–Bogdanov [77, 78, 79, 80, 81, 82, 83]). Such tasks would be intractable in the high-dimensional space generated by the full Navier–Stokes simulator, where one has to rely on linear (or weakly nonlinear) stability analysis of equilibria and on local theoretical normal-forms to trace bifurcation diagrams [83].
@@ 2 Methodology
The pipeline—rooted to the celebrated Equation-Free (EF) mutliscale numerical analysis framework [75]—for the systematic data-driven bifurcation and stability analysis of the Navier-Stokes equations consists of four main steps (see Fig. 2): (a) use manifold learning to “restrict/embed” the high-dimensional spatio-tempotal patterns into appropriate latent spaces, (b) learning surrogate ROMs in these latent spaces (here using Gaussian Regression Processes, i.e. GPR), (c) performing bifurcation and stability analysis in the latent spaces, (d) “lifting” to the high-dimensional space. A key distinction between our proposed framework and the traditional EF approach that was also coupled with POD for incompressible flows [84] is that, in the latter, one constructs local mappings and computes the quantities needed for bifurcation analysis via short bursts of microscopic simulations to bridge detailed simulations and emergent dynamics (captured by a few moment variables). In contrast, the proposed method is an extension to the traditional EF framework, in the sense that it learns a global model via machine learning from long-term, high-fidelity simulations spanning the entire parameter space. This global perspective enables the use of advanced nonlinear manifold-learning techniques, such as diffusion maps, and techniques for the solution of the pre-image problem, such as geometric harmonics or autoencoders, to construct maps between the high-dimensional and the latent space for complex dynamics like turbulence. As we have demonstrated in a series of works [57, 39, 85, 41, 61, 86], this framework can accurately reproduce the emergent bifurcation diagrams for simple one-dimensional PDEs and time series. Here, the framework is exploited to deal with the rich spatio-temporal complexity of the Navier–Stokes equations, representing a significant advance. We show for the first time that the proposed pipeline efficiently uncovers a low-dimensional representation that can accurately approximate the emergent dynamics via global surrogate models, whose governing equations closely resemble normal-forms of the correct dimension. This in turn allows the use of the state-of-the-art numerical bifurcation analysis toolkit, ensuring fidelity across a wide range of Reynolds numbers and the execution of tasks such as the continuation of limit cycles and their period, as well as their stability analysis. The latter task is done for the first time and showcases the potential of the framework to perform advanced bifurcation analysis. For the broader CFD community, the framework signals a promising new direction: the construction of fully data-driven normal-form ROMs, offering both high accuracy and efficiency, reducing computational expense by orders of magnitude. This will facilitate the execution of important tasks in CFD going beyond bifurcation analysis, such as real-time optimization and control.
@@
The representational fidelity of ROMs for accurately describing high-dimensional systems, like the Navier-Stokes equations (1a,1b, 1c), relies on the key assumption that the long-term dynamics evolve on a low-dimensional inertial manifold. This manifold is a finite-dimensional, attracting, and invariant subset [87, 88, 89] embedded in the full-order model’s infinite-dimensional phase space. In practice, ROMs are constructed by identifying approximate inertial manifolds (AIMs), which can be parameterized using a small number of dominant modes. A paradigmatic example is the celebrated Lorenz equations, derived as a three-mode reduction of the Rayleigh-Bénard convection equations, which retains the chaotic attractor characteristics of the full-order PDE. Data-driven AIM parameterizations can be constructed numerically using, for example, POD embeddings or nonlinear manifold learning embeddings, as recently shown [42], but also autoencoders [49], for highly nonlinear phenomena. Such data-driven algorithms complement analytical AIM theory by providing latent coordinates for the slow/long-term dynamics, enabling reduced modeling of complex dissipative systems from data. For a detailed discussion between data-driven ROMs and AIMs, see [49].
@@ 2.1 The four-stage proposed data-driven framework
Our aim, to first construct a non-intrusive dynamical ROM of minimal dimensions, reflecting a normal-form-ROM, by assuming a given set $X\in\mathbb{R}^{N\times m}$ of $M$ spatio-temporal snapshots $\{\mathbf{u}_{m}\}_{m=1,\ldots,M}$ that arise from the discretized Navier-Stokes equations, or from fluid flow experiments. Each snapshot at time $t_{m}$ contains the velocity variables across the spatial domain; i.e., $\mathbf{u}_{m}=[u(x_{i},y_{j},t_{m}),v(x_{i},y_{j},t_{m})]^{\top}\in\mathbb{R}^{2N_{g}}$, where $i=1,\ldots,N_{x}$, $j=1,\ldots,N_{y}$ and $N_{g}=N_{x}\times N_{y}$ is the number of grid points. Based on the above rationale, we furthermore assume that these points lie on a smooth $d$-dimensional manifold $\mathcal{M}\subset\mathbb{R}^{N}$, and we seek a data-driven mapping:
@@
$\hat{\boldsymbol{c}}\;=\;\arg\min_{\boldsymbol{c}}\Bigl{\|}\,\boldsymbol{a}_{*}\;-\;\varPhi\bigl{(}\varPsi(\boldsymbol{a}_{*};\boldsymbol{c})\bigr{)}\Bigr{\|}_{2},$
@@ 2.4 Gaussian Process regression
$\begin{bmatrix}\frac{d\mathbf{a}_{i}}{dt}\\ g_{i}(\bm{z}_{*})\end{bmatrix}\sim\mathcal{N}\left(\bm{0},\begin{bmatrix}\bm{K}_{i}(\bm{Z},\bm{Z}\mid\bm{\theta}_{i})+(\sigma_{i})^{2}\bm{I}_{M}&\bm{k}_{i}(\bm{Z},\bm{z}_{*}\mid\bm{\theta}_{i})\\ \bm{k}_{i}(\bm{z}_{*},\bm{Z}\mid\bm{\theta}_{i})&k_{i}(\bm{z}_{*},\bm{z}_{*}\mid\bm{\theta}_{i})\end{bmatrix}\right).$ (10)
@@
$-\log P\left(\tfrac{d\mathbf{a}_{i}}{dt}\mid\bm{Z},\bm{\theta}_{i}\right)=\frac{1}{2}\left(\tfrac{d\mathbf{a}_{i}}{dt}\right)^{\top}\bm{\Sigma}_{i}^{-1}\tfrac{d\mathbf{a}_{i}}{dt}+\frac{1}{2}\log|\bm{\Sigma}_{i}|+\frac{M}{2}\log 2\pi,$ (13)
@@
[…]

### v2

@@ Abstract
Inspired by the Equation-Free multiscale modelling approach, we show how the “embed-learn-lift” framework can be exploited for the construction of surrogate global “normal-forms” (reduced-order models of minimal dimensionality) of high-fidelity Navier-Stokes simulations, that can in turn be used for computationally efficient and numerically accurate bifurcation and stability analysis tasks, thus dealing with the presence of continuous symmetries. The framework consists of four main steps. In the first step, we apply manifold learning to discover the intrinsic dimension of the emergent latent space of the complex, high-dimensional Navier-Stokes spatio-temporal dynamics over the parameter space. In the second stage, we construct “normal-forms” like reduced-order models (ROMs)-in the sense that we capture the correct dimension for describing the emergent dynamics- on the latent space using Gaussian Process regression (GPR). Then, based on the constructed surrogate “normal-form” models, which dramatically reduce computational complexity, we exploit the toolkit of numerical bifurcation analysis theory to construct the bifurcation diagrams in the latent spaces, and perform systematic stability analysis. This allows, for example, the continuation of branches of limit cycles emanating from Andronov-Hopf bifurcations; such an analysis is intractable for high-dimensional systems such as the Navier-Stokes equations using standard Jacobian-full based computations. Finally, by solving the pre-image problem, we reconstruct the bifurcation diagrams in the original high-dimensional space. The effectiveness of the approach is demonstrated by focusing on two suitable test-bed flow configurations: the wake flow past an infinitely long circular cylinder and the planar sudden-expansion channel flow. These two flow systems exhibit Andronov-Hopf and pitchfork (symmetry-breaking) bifurcations by increasing the Reynolds number $Re$, respectively. The proposed methodology successfully identifies the appropriate dimensionality of the latent space of the Navier-Stokes equations—and based on them constructs ROMs referred to here as "surrogate normal-forms" using Gaussian Process regression (GPR)—that enable the accurate tracing of the bifurcating solutions and their stability analysis. This includes the accurate tracing of limit cycles, their period, and their stability characteristics, as determined through the computation of Floquet multipliers.
@@ 1 Introduction
On the other side, data-driven methods have been propelled in recent years by the exponential growth of machine learning (ML) and deep learning methodologies, which aim to construct surrogate ROMs from high-fidelity spatio-temporal simulations [16, 28, 29, 30]. The first step involves projecting the high-dimensional spatio-temporal flow dynamics onto a carefully chosen (low-dimensional) latent space. This in CFD is typically done via manifold learning techniques like Proper Orthogonal Decomposition (POD) [31, 32, 33, 34]. POD offers a direct, closed-form solution to the pre-image problem, i.e., the reconstruction of the CFD dynamics back in the original high-dimensional space. Nonlinear manifold learning algorithms, such as diffusion maps (DMs), have also been exploited  [35, 36, 37, 38, 39, 40, 41, 42] to deal with more complex dynamics, including chaotic flows [42]. In these approaches, in contrast to the straightforward closed-form mapping between full and reduced spaces provided by POD, the solution of the pre-image problem suffers from ill-posedness, and different techniques such as geometric harmonics and kernel ridge regression have been proposed (for a review, see [43, 44]). Machine learning-based architectures, such as autoencoders, have also been used to learn a set of coordinates that can parameterize the dynamics of the latent space [45, 46, 47, 48, 49]. On the part of the construction of surrogate ROMs via machine learning, various approaches have been proposed including Sparse identification of nonlinear dynamical systems (SINDy) [50, 51, 52, 53] coupled also with auto-encoders  [54, 55, 56], feedforward/recurrent neural networks [39, 40, 57, 58, 59, 60, 61, 62, 63], cluster-based networks [64], spectral submanifolds [65] and gaussian process regression models (GPR) [66, 67, 68, 69]. For example, SINDy has been used to model a range of fluid flows, including laminar and turbulent wakes [51], convective [52] and shear [53] flows, and it has been also applied in combination with auto-encoders to learn latent space coordinates before finding the dynamical system to evolve [54, 56]. Deep Neural Network-based Auto Encoders (DNN-AE) have been recently employed to identify reduced-order coordinates of a turbulent flow over a building-like structure on a nonlinear manifold [48]. Diffusion maps and neural networks have been leveraged to build predictive reduced-order models of a vertically falling liquid film flow in the case of sparse data [40]. A combination of $\beta$-variational autoencoders and transformers has been used to learn parsimonious and near-orthogonal reduced-order models for two-dimensional viscous flows in both periodic and chaotic regimes [70]. In [17, 71, 72], ROMs and bifurcation diagrams of equilibria are built from steady-state snapshots of Navier–Stokes simulations, targeting “stable bifurcating branches” using intrusive POD Galerkin projections with applications to triangular lid-driven cavity and sudden-expansion channel flows. This approach has also been used for the control of solutions on the steady-state unstable branch [73] and generalized in cases of fluid-structure interaction problems [74]. In [56], POD and Auto Encoders are combined with parametric SINDy to construct low-dimensional models from limited full-order data, enabling efficient bifurcation analysis of dynamical systems, with applications ranging from structural mechanics (clamped–clamped beam) to fluid dynamics (flow past a cylinder).
Here, building on previous efforts rooted in the celebrated Equation-Free (EF) multiscale framework [75, 76] introduced back in 2000’s, that is based on the on-demand construction of local models, we propose a four-stage full data-driven framework for the construction of surrogate global “normal-forms” via machine learning that can approximate not only qualitatively (as the traditional normal-form models) but also quantitatively, with high accuracy, the emergent dynamics of the Navier-Stokes equations, thereby enabling their use in numerical bifurcation and stability analysis, thus dealing also with the presence of continuous symmetries. In particular, in [76], we showed that a high-fidelity spherical-harmonics simulator of the Smoluchowski PDE for complex flows of liquid crystals can be exploited to carry out reduced, symmetry-factored bifurcation calculations. The only prerequisites of the EF approach are (i) that the full model is in principle reducible and (ii) a good choice of reduced observables that parametrize an attracting slow manifold; no explicit closure is required because short, appropriately initialized bursts of the full code provide on-demand estimates near that manifold. We also connected this timestepper approach to the Rowley–Marsden template method to dynamically factor out continuous symmetries (mesoscopic, e.g. liquid-crystal orientation, and macroscopic, e.g. translational/rotational symmetries in flows), making the method applicable to traveling/rotating solutions and to self-similar (scale-invariant) phenomena.
@@
Here, we extend this EF multiscale framework to discover global ROMs via machine learning. At this point, we note that we use the term “normal-forms” to emphasize the (main) challenge in complex systems modelling, that of learning the correct dimensionality of the latent dynamics across the parameter space, rather than reproducing the specific ROMs with minimal polynomial structures or local truncations as meant in local bifurcation theory. Consequently, our “normal-forms” for the Navier-Stokes equations represent ROMs with the correct dimensionality needed to approximate quantitatively the dynamics beyond the limited vicinity of the bifurcation points, which classical normal-forms cannot. In addition, machine-learning-trained surrogate models do not, by default, respect the symmetry properties of the high-fidelity simulator. Consequently, they can break symmetries, producing biased or spurious predictions, thus distorting the bifurcation structure or long-term dynamics. Here, to preserve those symmetries, we use explicit symmetry-aware functions as proposed in [77]. Based on the fully data-driven constructed surrogate models, we then exploit the power of state-of-the-art continuation toolboxes such as MATCONT [3, 78] and AUTO [2], to perform computationally efficient and highly accurate bifurcation and stability analysis of the Navier-Stokes equations. Hence, we dramatically reduce computational and memory demands, yet still allow all the state-of-the-art bifurcation and stability (continuation of solution branches, even of unstable steady-states, continuation of limit cycles and their exact period, branch switching, the calculation of eigenvalues and Floquet multipliers, continuation of codimension-two bifurcations, continuation of homoclinic and heteroclinic bifurcations, Bautin and Takens–Bogdanov [79, 80, 81, 82, 83, 84]). Such tasks would be intractable in the high-dimensional space generated by the full Navier–Stokes simulator, where one has to rely on linear (or weakly nonlinear) stability analysis of equilibria and on local theoretical normal-forms to trace bifurcation diagrams [84].
@@ 2 Methodology
The pipeline—rooted to the celebrated Equation-Free (EF) mutliscale numerical analysis framework [75]—for the systematic data-driven bifurcation and stability analysis of the Navier-Stokes equations consists of four main steps (see Fig. 1): (a) use manifold learning to “restrict/embed” the high-dimensional spatio-tempotal patterns into appropriate latent spaces, (b) learning surrogate ROMs in these latent spaces (here using Gaussian Regression Processes, i.e. GPR), (c) performing bifurcation and stability analysis in the latent spaces, (d) “lifting” to the high-dimensional space. A key distinction between our proposed framework and the traditional EF approach that was also coupled with POD for incompressible flows [85] is that, in the latter, one constructs local mappings and computes the quantities needed for bifurcation analysis via short bursts of microscopic simulations to bridge detailed simulations and emergent dynamics (captured by a few moment variables). In contrast, the proposed method is an extension to the traditional EF framework, in the sense that it learns a global model via machine learning from long-term, high-fidelity simulations spanning the entire parameter space. This global perspective enables the use of advanced nonlinear manifold-learning techniques, such as diffusion maps, and techniques for the solution of the pre-image problem, such as geometric harmonics or autoencoders, to construct maps between the high-dimensional and the latent space for complex dynamics like turbulence. As we have demonstrated in a series of works [41, 44, 58, 62, 76, 86], this framework can accurately reproduce the emergent bifurcation diagrams for simple one-dimensional PDEs and time series. Here, the framework is exploited to deal with the rich spatio-temporal complexity of the Navier–Stokes equations, representing a significant advance, thus extending the EF framework for the bifurcation analysis of high-fidelity simulators introduced in [75, 76] including simulators of complex fluids. Here, we show for the first time that the proposed pipeline efficiently uncovers a low-dimensional representation that can accurately approximate the emergent dynamics via global surrogate models, whose governing equations closely resemble normal-forms of the correct dimension. This in turn allows the use of the state-of-the-art numerical bifurcation analysis toolkit, ensuring fidelity across a wide range of Reynolds numbers and the execution of tasks such as the continuation of limit cycles and their period, as well as their stability analysis. The latter task is performed for the first time and showcases the potential of the framework to perform advanced bifurcation analysis. For the broader CFD community, the framework signals a promising new direction: the construction of fully data-driven normal-form ROMs, offering both high accuracy and efficiency, reducing computational expense by orders of magnitude. This will facilitate the execution of important tasks in CFD going beyond bifurcation analysis, such as real-time optimization and control.
@@
The representational fidelity of ROMs for accurately describing high-dimensional systems, like the Navier-Stokes equations (1a,1b, 1c), relies on the key assumption that long-term dynamics evolve on a low-dimensional inertial manifold. This manifold is a finite-dimensional, attracting, and invariant subset [87, 88, 89] embedded in the infinite-dimensional phase space of the full-order model. In practice, ROMs are constructed by identifying approximate inertial manifolds (AIMs), which can be parameterized using a small number of dominant modes. A paradigmatic example are the celebrated Lorenz equations, derived as a three-mode reduction of the Rayleigh-Bénard convection equations, which retains the chaotic attractor characteristics of the full-order PDE. Data-driven AIM parameterizations can be constructed numerically using, for example, POD embeddings or nonlinear manifold learning embeddings, as recently shown [42], but also autoencoders [49], for highly nonlinear phenomena. Such data-driven algorithms complement analytical AIM theory by providing latent coordinates for the slow/long-term dynamics, enabling reduced modeling of complex dissipative systems from data. For a detailed discussion between data-driven ROMs and AIMs, see [49].
@@ 2.1 The four-stage proposed data-driven framework
Our aim, to first construct a non-intrusive dynamical ROM of minimal dimensions, reflecting a normal-form ROM, by assuming a given set $X\in\mathbb{R}^{N\times m}$ of $M$ spatio-temporal snapshots $\{\mathbf{u}_{m}\}_{m=1,\ldots,M}$ that arise from the discretized Navier-Stokes equations, or from fluid flow experiments. Each snapshot at time $t_{m}$ contains the velocity variables across the spatial domain; i.e., $\mathbf{u}_{m}=[u(x_{i},y_{j},t_{m}),v(x_{i},y_{j},t_{m})]^{\top}\in\mathbb{R}^{2N_{g}}$, where $i=1,\ldots,N_{x}$, $j=1,\ldots,N_{y}$ and $N_{g}=N_{x}\times N_{y}$ is the number of grid points. Based on the above rationale, we furthermore assume that these points lie on a smooth $d$-dimensional manifold $\mathcal{M}\subset\mathbb{R}^{N}$, and we seek a data-driven mapping:
@@
$\hat{\boldsymbol{c}}\;=\;\arg\min_{\boldsymbol{c}}\Bigl{\|}\,\boldsymbol{a}_{*}\;-\;\varPhi\bigl{(}\varPsi(\boldsymbol{a}_{*};\boldsymbol{c})\bigr{)}\Bigr{\|}_{2},$ (5)
@@ 2.4 Gaussian Process regression
[…]
@@
$\log P\left(\tfrac{d\mathbf{a}_{i}}{dt}\mid\bm{Z},\bm{\theta}_{i}\right)=\frac{1}{2}\left(\tfrac{d\mathbf{a}_{i}}{dt}\right)^{\top}\bm{\Sigma}_{i}^{-1}\tfrac{d\mathbf{a}_{i}}{dt}+\frac{1}{2}\log|\bm{\Sigma}_{i}|+\frac{M}{2}\log 2\pi,$ (14)
@@
$\left\{\begin{aligned} \dfrac{da_{1}}{dt}&=g_{1}(a_{1},\dots,a_{d},Re),\\ &\vdots\\ \dfrac{da_{d}}{dt}&=g_{d}(a_{1},\dots,a_{d},Re).\end{aligned}\right$. (16)